OpenStudy (anonymous):

let X be a binomial variable with parameters n and p .can someone help with this question

OpenStudy (anonymous):

man this must be a hell of a question

OpenStudy (anonymous):

let X be a binpmial variable with parameters n and p , such that np=\[\lambda \.show that for fixed k ,as \[ n \rightarrow \infty ,with \lambda fixed ,\] \[ p(X=k)\rightarrow (1/k !)\lambda^{k}e^{-\lambda}\]

OpenStudy (anonymous):

this is the poisson distribution. proof will be in any text

OpenStudy (anonymous):

but i dont get the answer

OpenStudy (anonymous):

i can grab a text and start copying, but really it is in any book under the heading something like "the poisson as an approximation to the binomial" are you reading from a book? as i recall it is not hard

OpenStudy (anonymous):

if so let me know we can go through it step by step

OpenStudy (anonymous):

can u solve it to me i'm stuck

OpenStudy (anonymous):

i have in front of me three proofs. all are basically the same and require one somewhat tricky piece of algebra. are you looking at a proof?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but it does not have step they just says is a poisson then they give that answer i've showed you

OpenStudy (anonymous):

ok so we know binomial is \[P(X=k)=\dbinom{n}{k}p^k(1-p)^{n-k}\] and now \(\lambda =np \implies p=\frac{\lambda}{n}\) so step one is to put \[P(X=k)=\dbinom{n}{k}(\frac{\lambda}{n})^k(1-\frac{\lambda}{n})^{n-k}\]

OpenStudy (anonymous):

actually maybe that is the last step. here is a more straightforward solution we have \[\dbinom{n}{k}p^k(1-p)^{n-k}=\frac{n(n-1)...(n-k+1)}{k!}p^k(1-p)^{n-k}\] and keeping in mind that as \(n\to \infty\) we have \(np -\to \lambda\) we can put \[\lambda^k=(np)^k=n^kp^k\] so replace and divide each term in the numerator by \(n^k\) to get \[\frac{(1-\frac{1}{n})(1-\frac{2}{n})...(1-\frac{(k-1)}{n})}{k!}(np)^k(1-p)^{n-k}\] also in the limit \(p\to \frac{\lambda}{n}\) so we can write \[\lim_{n\to \infty}\frac{(1-\frac{1}{n})(1-\frac{2}{n})...(1-\frac{(k-1)}{n})}{k!}(np)^k(1-p)^{n-k}\] \[=\lim_{n\to \infty}\frac{(1-\frac{1}{n})(1-\frac{2}{n})...(1-\frac{(k-1)}{n})}{k!}(\lambda)^k(1-\frac{\lambda}{n})^{n-k}\] \[=\lim_{n\to \infty}\frac{\lambda^k}{k!}(1-\frac{\lambda}{n})^{n-k}\] because the limit in that ugly numerator is 1

OpenStudy (anonymous):

last step is to compute \[\lim_{n\to \infty}(1-\frac{\lambda}{n})^{n-k}\] but k is fixed and only \(n\to \infty\) so this is the same as \[\lim_{n\to \infty}(1-\frac{\lambda}{n})^n\] which is certainly \[e^{-\lambda}\]

OpenStudy (anonymous):

i will be proficient \(\LaTeX\) soon

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

yw